OpenStudy (anonymous):

can somebody help me graph this?? 4x + y = 5 3x – 7y = 27

OpenStudy (anonymous):

Need a graph paper to do so! But it's not available in here :S

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=plot+graphs+4x+%2B+y+%3D+5 this for one of the equations Hope it Helps

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

4x+y=5*3x-7y=27 4 x+y = 15 x-7 y = 27 x = 216/43, y = 297/43 Here!